# Householder Pseudo-Rotation: A Novel Approach to Activation Editing in LLMs with Direction-Magnitude Perspective

### v1

@@ Abstract
Activation Editing, which involves directly editting the internal representations of large language models (LLMs) to alter their behaviors and achieve desired properties, has emerged as a promising area of research. Existing works primarily treat LLMs’ activations as points in space and modify them by adding steering vectors. However, this approach is limited in its ability to achieve greater performance improvement while maintaining the necessary consistency of activation magnitudes. To overcome these issues, we propose a novel editing method that views activations in terms of their directions and magnitudes. Our method, named Householder Pseudo-Rotation (HPR), mimics the rotation transformation, thus preserving activation norms and resulting in an improved performance on various safety benchmarks.
@@ 1 Introduction
In an experiment with the activation space, we discover an important property that are maintained by powerful LLMs: activations within the same layer tend to have roughly the same vector norm. We refer to this as the Magnitude Consistency property, i.e., Section 4.3. This observation highlights a key limitation of the points-in-space view, where the steering vector approach cannot simultaneously maintain activation magnitude consistency and effective activation editing to achieve greater performance improvement for desired behaviors for LLMs. If the scaling factor is too large, the additive edit might drastically alter the activation norms in each layer, violating the norm consistency property of LLMs. In extreme cases, this change can lead to the generation of complete gibberish, undermining the desired behaviors of the LLM’s responses. Conversely, if the scaling factor is set too low to preserve the activation norms, the steering vector may have limited abilities to shift an activation toward new behavior, thus also hindering editing performance for desired behaviors. Moreover, the steering vector approach does not align with the commonly used cosine similarity metric, which emphasizes directional alignment between vectors rather than their absolute positions.
@@ 2.1 Problem Statement
Let =$\mathcal{M}=\{\mathcal{M}^{(l)}|0\leq l<L\}$ be a $L$-layers pre-trained LLM whose behavior needs to be altered. Assume that the outputs of $\mathcal{M}$ exhibit either of the two contrasting qualities: a positive behavior $\mathbf{p}$ or a negative behavior $\mathbf{n}$. For instance, $\mathbf{p}$ can be truthfulness and $\mathbf{n}$ is untruthfulness. We denote:
@@
When the label of the output, i.e. positive or negative, is unknown, we refer to its length as $S^{y}$.
@@ 3.3 Computing the Final Activation
$\hat{\sigma}$ is rounded to the nearest integer, 00 or $1$ to be specific, and predict whether the given activation $a$ is positive or negative. If $a$ is predicted as a negative activation, we edit it by first reflecting $a$ about the separating hyperplane $\theta_{probe}$ to obtain the reflected vector $\dot{a}$ in the positive region. Afterward, we calculate a new activation by rotating $a$ within the $2$-D plan formed by $a$ and $\dot{a}$ by an angle of $\gamma_{1}$ radians. The resulting vector $\hat{a}$ will serve as our predicted positive activation for $a$.
@@ 4.1 Experimental Setup
Hyperparameters: In our model, the linear probe is a vector of the same dimensions as the LLMs’ hidden dimensions. The angle prediction module is a feedforward neural network with 4 layers and one output unit. We train each module for 5 epochs with batch size $16$, AdamW optimizer (Loshchilov and Hutter, 2019), learning rate $5\times 10^{-4}$, cosine learning rate scheduler and warmup. For editing, we apply HPR to the top $k=5$ layers with the highest probe accuracy. Appendix C presents model performance with different values of $k$. We also provide a reproducibility checklist in Appendix LABEL:app:repo.
@@ 4.2 Results
Ablation Study: The last row in Table 1 further shows the performance of HPR when the angle prediction module is excluded from the full model. As can be seen, this exclusion leads to significant performance drops across all base LLMs for HPR, thereby justifying the importance of angle prediction to adjust reflected activations for our model. We also note that the linear probe module cannot be removed from HPR for ablation study as it is essential for finding the positive-negative separating hyperplane and rotating plane in our model. Finally, the superior performance of HPR for different LLMs confirms the advantages of our assumption on the shared $2$-D plane of $a$, $\dot{a}$, and $\hat{a}$.
@@
BBQ, SEQ, and Toxigen: To further illustrate the effectiveness of HPR in eliciting desirable behavior, Table 2 shows HPR’s performance on the BBQ, SEQ, and Toxigen datasets. These datasets evaluate the abilities of LLMs to generate unbiased (BBQ), ethically acceptable (SEQ), and non-toxic (Toxigen) responses. Across various base LLMs, incorporating HPR can significantly enhance performance on all these datasets. These results highlight the benefits of HPR in improving important safety criteria for LLMs, leading to unbiased, ethical, and non-toxic responses for responsible models.
@@ 4.3 Analysis of Activation Space
What is the implication of such slight norm divergence from base LLMs for ITI? In Table 3, we present the behavior shift rates of ITI15 and ITI200 compared to the original Llama2-7B-Chat model in TruthfulQA. Specifically, we show how often each editing method can flip the LLM’s predictions of examples from true to false and vice versa. From the table, we observe that the slight divergence of activation norms in ITI15 results in a more limited ability to change the base model’s behavior, with a behavior shift rate of only $8.56\%$ compared to $34.23\%$ for ITI200. As the behavior shift rate is the upper bound of the overall performance improvement in TruthfulQA for ITI, this limited ability to alter LLM behavior will hinder further improvement with a small scaling factor in ITI.
@@ 6 Conclusion
Our work utilize open-source LLMs, i.e., Llama2-7B-Chat (Touvron et al., 2023b), Mistral-7B-Instruct (Jiang et al., 2023), and Llama3-8B-Instruct (AI@Meta, 2024), as the base models, thus potentially inheriting their inherent societal issues like bias, hallucination, privacy, etc. Simultaneously, we propose a novel activation editing method aiming at altering LLMs’ behaviour for the better, contributing to the on-going efforts to advance LLM safety. As activation and model editing for LLMs has been studied in recent published work Li et al. (2023b); Liu et al. (2021); Ilharco et al. (2023), we do not believe our work poses greater societal risks than such studies and open-source LLMs. Finally, we confirm that we follow all the ethical guideline from ACL ARR to the best of our knowledge when performing this research.
@@ Appendix B Training Efficiency
During the training phase, we use $a_{i,j}^{(l),\mathbf{p}}$ / $a_{i,j}^{(l),\mathbf{n}}$ pairs to form the inputs and labels for the linear probe and angle prediction modules in each layer. Generally, these are computed by passing training data samples $x\|y_{i}^{\mathbf{p}}$ and $x\|y_{i}^{\mathbf{n}}$ through the model $\mathcal{M}$ and record the activations at each layer and token position. However, since our method does not update the parameters of $\mathcal{M}$, its activation vectors can be treated as constants. Thus, before training we pre-compute all activations on the training data to make a dataset of $a_{i,j}^{(l),\mathbf{p}}$ / $a_{i,j}^{(l),\mathbf{n}}$ pairs for each layer. These can then be used to train the linear probe and angle prediction modules independently of the base model. In this way, the base LLM does not need to be loaded into GPU RAM, saving more space for training the HPR modules.

### v2

@@ Abstract
Activation Editing, which involves directly editing the internal representations of large language models (LLMs) to alter their behaviors and achieve desired properties, has emerged as a promising area of research. Existing works primarily treat LLMs’ activations as points in space and modify them by adding steering vectors. However, this approach is limited in its ability to achieve greater performance improvement while maintaining the necessary consistency of activation magnitudes. To overcome these issues, we propose a novel editing method that views activations in terms of their directions and magnitudes. Our method, named Householder Pseudo-Rotation (HPR), mimics the rotation transformation, thus preserving activation norms and resulting in an improved performance on various safety benchmarks.
@@ 1 Introduction
In an experiment with the activation space, we discover an important property that is maintained by powerful LLMs: activations within the same layer tend to have roughly the same vector norm. We refer to this as the Magnitude Consistency property, i.e., Section 4.3. This observation highlights a key limitation of the points-in-space view, where the steering vector approach cannot simultaneously maintain activation magnitude consistency and effectively edit activation to achieve greater performance improvement for desired behaviors for LLMs. If the scaling factor is too large, the additive edit might drastically alter the activation norms in each layer, violating the norm consistency property of LLMs. In extreme cases, this change can lead to the generation of complete gibberish, undermining the desired behaviors of the LLM’s responses. Conversely, if the scaling factor is set too low to preserve the activation norms, the steering vector may have limited abilities to shift an activation toward new behavior, thus also hindering editing performance for desired behaviors. Moreover, the steering vector approach does not align with the commonly used cosine similarity metric, which emphasizes directional alignment between vectors rather than their absolute positions.
@@ 2.1 Problem Statement
Let =$\mathcal{M}=\{\mathcal{M}^{(l)}|0\leq l<L\}$ be a $L$-layers pre-trained LLM whose behavior needs to be altered. Assume that the outputs of $\mathcal{M}$ exhibit either of the two contrasting qualities: a positive behavior $\mathbf{p}$ or a negative behavior $\mathbf{n}$. For instance, $\mathbf{p}$ and $\mathbf{n}$ can be truthfulness and untruthfulness. We denote:
@@
Here, $i$ is the sample index in the dataset, and $j$ is the token index in a sample. When the label of the output, i.e. positive or negative, is unknown, we refer to its length as $S^{y}$.
@@ 3.3 Computing the Final Activation
$\hat{\sigma}$ is rounded to the nearest integer, 00 or $1$ to be specific, and predicts whether the given activation $a$ is positive or negative. If $a$ is predicted as a negative activation, we edit it by first reflecting $a$ about the separating hyperplane $\theta_{probe}$ to obtain the reflected vector $\dot{a}$ in the positive region. Afterward, we calculate a new activation by rotating $a$ within the $2$-D plan formed by $a$ and $\dot{a}$ by an angle of $\gamma_{1}$ radians. The resulting vector $\hat{a}$ will serve as our predicted positive activation for $a$.
@@ 4.1 Experimental Setup
Hyperparameters: In our model, the linear probe is a vector of the same dimensions as the LLMs’ hidden dimensions. The angle prediction module is a feedforward neural network with 4 layers and one output unit. We train each module for 5 epochs with batch size $16$, AdamW optimizer (Loshchilov and Hutter, 2019), learning rate $5\times 10^{-4}$, cosine learning rate scheduler and warmup. For editing, we apply HPR to the top $k=5$ layers with the highest probe accuracy. Appendix C presents model performance with different values of $k$.
@@ 4.2 Results
Ablation Study: The last row in Table 1 further shows the performance of HPR when the angle prediction module is excluded from the full model. In other words, the editing function now only reflects negative activation vectors about the separating hyperplane defined by the linear probe. As can be seen, this exclusion leads to significant performance drops across all base LLMs for HPR, suggesting that simply having the activations landed on the positive region is not enough to make an effective edit. Thereby, it justifies the importance of angle prediction to adjust reflected activations for our model. We also note that the linear probe module cannot be removed from HPR for ablation study as it is essential for finding the positive-negative separating hyperplane and rotating plane in our model. Finally, the superior performance of HPR for different LLMs confirms the advantages of our assumption on the shared $2$-D plane of $a$, $\dot{a}$, and $\hat{a}$.
@@
BBQ, SEQ, and Toxigen: To further illustrate the effectiveness of HPR in eliciting desirable behavior, Table 2 shows HPR’s performance on the BBQ, SEQ, and Toxigen datasets. These datasets evaluate the abilities of LLMs to generate unbiased (BBQ), ethically acceptable (SEQ), and non-toxic (Toxigen) responses. Across various base LLMs, incorporating HPR can significantly enhance performance on all of these datasets. These results highlight the benefits of HPR in improving important safety criteria for LLMs, leading to unbiased, ethical, and non-toxic responses for responsible models.
@@ 4.3 Analysis of Activation Space
What is the implication of such slight norm divergence from base LLMs for ITI? In Table 3, we present the behavior shift rates of ITI15 and ITI200 compared to the original Llama2-7B-Chat model on TruthfulQA. Specifically, we show how often each editing method can flip the LLM’s predictions of examples from true to false and vice versa. From the table, we observe that the slight divergence of activation norms in ITI15 results in a more limited ability to change the base model’s behavior, with a behavior shift rate of only $8.56\%$ compared to $34.23\%$ for ITI200. As the behavior shift rate is the upper bound of the overall performance improvement on TruthfulQA for ITI, this limited ability to alter LLM behavior will hinder further improvement with a small scaling factor in ITI.
@@
4.4 Impact on Generation Capability
Generation Quality: To assess the impact of the proposed method on generation capability, we perform evaluations in the open-ended generation setting of TruthfulQA with LLaMA models. For this evaluation, we employ BLEU accuracy, which is calculated as the ratio of generated responses having BLEU scores with their respective correct (positive) references higher than that with the incorrect (negative) references, as described in Lin et al. (2022). We use the popular implementation in Gao et al. (2023) and the same data split as in Section 4.1. The results are presented in Table 4, where ITI15 and ITI50 refer to the ITI method with a scaling factor $\alpha=15$ and $\alpha=50$ respectively.
Backbone Method BLEU Acc
Llama2-7B-Chat Base 38.39
ITI15 41.56
ITI50 34.96
HPR 42.30
Llama3-8B-Instruct Base 43.28
ITI15 41.32
ITI50 41.08
HPR 44.74
Table 4: Automated evaluation of TruthfulQA open-ended generation task.
It is clear from the table that the proposed HPR method also outperforms different baselines significantly on generation-based evaluation for TruthfulQA.
In addition, we evaluate the fluency of the models’ generated responses. Table 5 shows the perplexity scores and bits per byte (smaller is better) on the test set of WikiText-2 (Merity et al., 2017).
[…]
Llama2 Base 13.7077 1.6316 0.7063
ITI15 14.2038 1.6425 0.7158
ITI50 133.7374 2.4982 1.3209
HPR 13.7206 1.6319 0.7066
Llama3 Base 11.9524 1.5903 0.6693
ITI15 17.0515 1.6996 0.7652
ITI50 4303.4616 4.7813 2.2574
HPR 11.9558 1.5904 0.6694
Table 5: Perplexity and bits-per-byte on WikiText-2 test set.
As can be seen, our method HPR does not need to sacrifice the models’ fluency to achieve effective editing for desirable model behavior, unlike the Steering Vector methods such as ITI. Notably, when the scaling factor of ITI is high (i.e., $50$), the perplexity scores become extremely large, leading to gibberish responses.
Llama2-7B Llama3-8B Mistral-7B
Base 2.33 2.08 2.33
LoRA 1.30 1.24 1.23
ITI 2.06 1.87 -
HPR1 2.18 1.99 2.25
HPR5 1.98 1.80 1.98
HPR10 1.73 1.63 1.76
Table 6: Inference speed in samples per second (larger is better).
Inference Speed: Table 6 compares HPR with the base model, ITI, and LoRA adapter in terms of inference speed. For HPR, we report the speed for its three variants corresponding to the number of edited layers. We see that with only one edited layer, the inference speed is slower than the base model but faster than LoRA and ITI. HPR slows down when we choose more layers to edit, which is natural. Importantly, we do not see significant speed reduction due to the introduction of our activation editing method, thus suggesting its potential applications in different scenarios.
In fact, efficiency is a key motivation for the design of our method. Our method does not perform direct rotation for each activation in the models as it will be very expensive. Instead, we find a common hyperplane to separate the negative and positive regions of the activations, and we use this hyperplane to efficiently find the rotating direction for all activations in the layer (i.e., via the reflections). As finding rotating directions is the most expensive part for a rotation operation, using a common hyperplane for all activations significantly reduces our computation costs for editing. Finally, to compute the rotating angles, our method employs a regression model, which is very efficient and can be applied for each activation to improve the performance of our method.
@@ 6 Conclusion
Our work utilize open-source LLMs, i.e., Llama2-7B-Chat (Touvron et al., 2023b), Mistral-7B-Instruct (Jiang et al., 2023), and Llama3-8B-Instruct (AI@Meta, 2024), as the base models, thus potentially inheriting their inherent societal issues like bias, hallucination, privacy, etc. Simultaneously, we propose a novel activation editing method aiming at altering LLMs’ behaviour for the better, contributing to the on-going efforts to advance LLM safety. As activation and model editing for LLMs has been studied in recent published work Li et al. (2023b); Liu et al. (2021); Ilharco et al. (2023), we do not believe our work poses greater societal risks than such studies and open-source LLMs. Finally, we confirm that we follow all the ethical guideline from ACL to the best of our knowledge when performing this research.
Acknowledgements
This research has been supported by the Army Research Office (ARO) grant W911NF-21-1-0112, the NSF grant CNS-1747798 to the IUCRC Center for Big Learning, and the NSF grant # 2239570. This research is also supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract 2022-22072200003. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government.
@@ Appendix B Training efficiency
During the training phase, we use $a_{i,j}^{(l),\mathbf{p}}$ / $a_{i,j}^{(l),\mathbf{n}}$ pairs to form the inputs and labels for the linear probe and angle prediction modules in each layer. Generally, these are computed by passing training data samples $x_{i}\|y_{i}^{\mathbf{p}}$ and $x_{i}\|y_{i}^{\mathbf{n}}$ through the model $\mathcal{M}$ and record the activations at each layer and token position. However, since our method does not update the parameters of $\mathcal{M}$, its activation vectors can be treated as constants. Thus, before training we pre-compute all activations on the training data to make a dataset of $a_{i,j}^{(l),\mathbf{p}}$ / $a_{i,j}^{(l),\mathbf{n}}$ pairs for each layer. These can then be used to train the linear probe and angle prediction modules independently of the base model. In this way, the base LLM does not need to be loaded into GPU RAM, saving more space for training the HPR modules.